# ME-Switch: A Memory-Efficient Expert Switching Framework for Large Language Models

### v1

@@ 4.1 Salient-Aware Delta Compression
4.2 Model-level Routing
In this section, we present a method to determine the appropriate model for a given user query. Consider a set of LLMs represented as =$\mathcal{F}=\{f_{1},f_{2},\cdots,f_{M}\}$, where $M$ denotes the number of models. Given a user query $q$, we aim to find the most suitable LLM by solving the following problem $\arg\max_{f\in\mathcal{F}}P(q,f(q))$, where $P$ is a function that measures the quality or performance of the LLM response. To simplify the routing process, we assume that each LLM in $\mathcal{F}$ specializes in distinct domains such as code generation or mathematical problem solving, This setup allows us to treat the routing challenge as a multiple-choice question-answering task, where each option corresponds to a specific domain, thereby transforming the problem into a domain classification problem. Note that dialogue LLMs like Qwen1.5-1.8B-Chat [8] exhibit capabilities in following instructions, which inspires us to utilize a small pre-trained LLM as a model-level router. As illustrated in Figure 3, we first prompt the router with the user’s question using a template designed to elicit domain classification. For the prompt template, please refer to Section A of the supplementary material. Based on the router’s response, we then dynamically load the corresponding compressed delta weights for the selected domain-specific model, such as a mathematical model, to generate outputs.
@@ 4.2 Model-level Routing
Discussion with Mixture of Experts (MoE). Besides model-level routing, another approach to handle diverse user queries efficiently is to construct a mixture-of-experts using a set of pre-trained models. This can be achieved by integrating the feedforward layers from all pre-trained LLMs into a single MoE module at each attention-FFN block, and merging other layers, such as self-attention layers, by simply averaging their weights [40]. An additional gate network is introduced for each MoE module to perform token-level routing. However, this approach requires extensive fine-tuning of the entire network parameters and the gating network, as there is a significant gap between MoE experts and pre-trained LLMs. Notably, each expert in an MoE tends to become a generalist across all domains due to the load balancing loss, which encourages an even distribution of the workload among experts [16, 23]. This contrasts with pre-trained LLMs, which are typically specialists within their specific domains.
Model size reduction analysis. Let $\Psi$ be the model size of an FP16 pre-trained model. When storing $M$ FP16 models, the total model size is $M\Psi$. In contrast, using our method, we store a single base model, a model-level router and $M$ compressed delta models. The total model size is $\Psi+M\tilde{\Psi}+\Phi$, where $\tilde{\Psi}$ represents the size of the compressed delta model and $\Phi$ denotes the storage requirement for the router. Therefore, the compression ratio can be computed by $M\Psi/(\Psi+M\tilde{\Psi}+\Phi)$.
@@ 5 Experiments
Candidate LLMs. We apply our ME-Switch to two model families, Mistral-7B [22] and LLaMA-2-13B [45]. For the Mistral family, we include Dolphin-2.2.1-Mistral-7B [3] as the instruction expert, Speechless-Code-Mistral-7B [5] as the code expert, and MetaMath-Mistral-7B [52] as the math expert. For the LLaMA-2-13B family, LLaMA-2-13B-Chat [45] serves as the instruction expert, MetaMath-13B [52] as the math expert, and LLaMA2-Chinese-13B-Chat [4] as the Chinese expert. The above models are fine-tuned based on pre-trained backbones. We use Qwen1.5-1.8B-Chat [8] as the model-level router.
Training and testing datasets. We collect a diverse set of instruction samples from various open-source datasets, including Alpaca [42] for the instruction domain, MetaMathQA [52] for the mathematics domain, Code-74k-ShareGPT [2] for the code domain, and Chinese Alpaca [1] for the Chinese domain. To evaluate the performance of the resulting LLMs, we report accuracy on several benchmarks across different domains: MMLU [17] for the instruction, GSM8K [11] and MATH [18] for the mathematics, HumanEval [10] and MBPP [7] for the code, and C-Eval [20] and C-MMLU [25] for the Chinese. We use the WizardCoder toolbox to evaluate on HumanEval and MBPP, and the OpenCompass toolbox [12] to evaluate the other datasets.
Implementation details. For salient-aware delta compression, we construct a calibration set from each domain-specific dataset and use these sets to compress the corresponding delta weights for each domain. Each calibration set consists of 1600 randomly sampled sequences, each with a length of 128 tokens. The bitwidth $b$ and the number of FP16 input channels $k$ are set to 2 and 8, respectively. We employ the AdamW optimizer [30] with a learning rate of $10^{-5}$ for 1 epoch, with a mini-batch size of 4. Delta weights compression experiments for the Mistral family are conducted on two NVIDIA A100 80G GPUs, while for the LLaMA-2-13B model, we use four NVIDIA A100 80G GPUs.
@@
ME-Switch 44.25 59.65 63.18 56.64 54.57 69.98 13.14 41.56 40.46 39.75 40.11
Table 2: Effect of salient-aware delta compression for Speechless-Code-Mistral-7B and MetaMath-Mistral-7B. $\tilde{\Delta}$ denotes the compressed delta weights.
Model Bit # FP Channels Model Size (GB) Code Generation (%) $\uparrow$
HumanEval MBPP Avg.
Speechless-Code-Mistral-7B 16 0 13.48 51.20 60.40 55.80
@@
Speechless-Code-Mistral-7B $\tilde{\Delta}$ 2 0 2.11 47.00 59.10 53.05
Speechless-Code-Mistral-7B $\tilde{\Delta}$ 2 8 2.13 51.80 60.70 56.25
Model Bit # FP Channels Model Size (GB) Mathematical Reasoning (%) $\uparrow$
GSM8K Math Avg.
MetaMath-Mistral-7B 16 0 13.48 73.92 20.62 47.27
@@
MetaMath-Mistral-7B $\tilde{\Delta}$ 2 0 2.11 73.31 20.44 46.88
MetaMath-Mistral-7B $\tilde{\Delta}$ 2 8 2.13 73.62 20.48 47.05
5.1 Main Results
To evaluate the efficacy of our proposed model, we apply ME-Switch to the Mistral-7B and LLaMA-13B model families. The experimental results, detailed in Table 1, demonstrate that ME-Switch, even with extremely compressed delta weights, achieves performance comparable to that of the respective unquantized expert models across various downstream tasks. For the Mistral-7B family, on MMLU, ME-Switch trails the instruction expert by only 0.1% and lags behind the math expert by just 0.4% in mathematical reasoning tasks. Notably, ME-Switch consistently outperforms the code expert in code generation tasks. The good performance of ME-Switch can be attributed to our accurate selection of salient delta weight channels, efficient distillation for step size learning, and precise routing for user queries. Furthermore, larger models typically contain a higher proportion of redundancy. Consequently, our ME-Switch exhibits even better performance on the LLaMA-13B family. For example, ME-Switch achieves an average accuracy improvement of 0.39% over the Chinese expert in the Chinese domain.
@@ 5.2 Ablation Studies
Effect of salient-aware delta compression. To validate the effect of our proposed salient-aware delta compression, we compress the delta weights of Speechless-Code-Mistral-7B and MetaMath-Mistral-7B with different bitwidths and evaluate their performance on code generation and mathematical reasoning tasks, respectively. The results presented in Table 2, indicate that quantizing delta weights to only 1-bit leads to a significant performance drop. Increasing the bitwidth to 2 improves performance, but it still lags behind that of the full-precision model. Notably, despite maintaining a minimal number of FP16 channels, the model size of the mixed-precision model (2.13 GB) is nearly unchanged compared to the fully quantized model (2.11 GB). However, introducing a small number of FP channels significantly improves performance. For example, our compressed Speechless-Code-Mistral-7B, with a model size reduced by 6.33$\times$, outperforms the full-precision counterpart by 0.45% in average accuracy on code generation tasks. This underscores the capability of salient-aware delta compression to minimize model size while preserving comprehensive model performance.
Table 3: Performance comparisons with other weight-only quantization methods.
Method FP 1-bit AWQ Random Wanda Magnitude Ours
HumanEval 51.20 47.00 47.00 46.30 48.20 48.20 47.60
MBPP 60.40 58.40 58.10 58.60 58.60 58.90 60.70
Avg. 55.80 52.70 52.55 52.45 53.40 53.55 54.15
@@
Performance comparisons with other weight-only quantization methods. To demonstrate the promising performance of our salient-aware delta compression, we also include the following weight-only quantization methods: AWQ: we use AWQ [26] to rescale input channels of delta weights before quantization to mitigate quantization errors. Random: using our salient-aware delta compression, we randomly select some input channels from the delta weights as important channels. Wanda: leveraging our salient-aware delta compression, we select important input channels of delta weights using the pruning metric from Wanda [41]. Magnitude: within our salient-aware delta compression framework, we select sensitive input channels of delta weights based on their weight magnitude, following the method proposed by [13]. For fair comparisons, we apply all methods to compress the delta weights of Speechless-Code-Mistral-7B to 1-bit. All mixed-precision methods retain 32 FP16 channels. We also include a 1-bit baseline for comparison. The results are shown in Table 3. From the results, we observe that AWQ achieves comparable performance to the 1-bit baseline, highlighting the limitations of rescaling in extremely low-bitwidth quantization. In contrast, keeping the salient delta input channels performs favourably against the rescaling input channel counterpart. Moreover, our salient channel selection demonstrates superior performance than Random, Wanda, and Magnitude metrics. For example, our salient-aware delta compression outperforms Wanda by 0.75% on the average accuracy, showcasing its effectiveness in selecting the salient delta weights.
@@
Model size reduction analysis. To investigate the model size reduction as discussed in Section 4.2, we compare the total storage requirements of full-precision models with those of our compressed models for the Mistral-7B and LLaMA-13B families across varying model counts, as shown in Figures 5 and 7. As the number of models increases, the compression ratios improve substantially. For instance, with nine models, our method achieves a 3.36$\times$ reduction compared to full-precision models. These savings become even more pronounced with larger model sizes, reaching up to a 3.63$\times$ reduction for LLaMA-13B family.
Latency analysis. To assess the latency improvements from delta weights compression, we measured the end-to-end decoding latency of the Mistral-7B model with an input sequence length of 128 on a single NVIDIA A100. Decoding latency is critical, as it typically dominates processing time in LLM operations [26, 29]. Our efficient Triton kernel, which enables batched matrix multiplication between multiple compressed weight matrices and high-precision input activations, is compared against the conventional approach of individually processing multiple models. Results depicted in Figure 7 illustrate that while our method may perform slightly slower than the naive approach for a small number of models, it provides lower latency as the number of models is greater than 4. More importantly, unlike the naive approach, our method is able to simultaneously serve 16 models on GPUs without running into out-of-memory (OOM) issues, demonstrating better scalability and efficiency in high-load scenarios.
@@ 6 Conclusion and Future Work
Broader Impacts. The substantial storage savings and competitive accuracy highlight ME-Switch’s potential for scalable and cost-effective AI deployments. The improved efficiency and scalability of ME-Switch could lead to the proliferation of LLMs, making it easier to deploy these models in various applications, including those that might be used maliciously. The widespread availability of powerful LLMs increases the risk of misuse in generating harmful content, spreading misinformation, and exacerbating issues related to bias and discrimination inherent in the training data.
Supplemental Material

### v2

@@ 4.2 Model-level Routing
4.3 Discussions
Comparisons with routing in Mixture of Experts (MoE). Besides model-level routing, another approach to handle diverse user queries efficiently is to construct a MoE using a set of pre-trained expert models. This can be achieved by integrating the feedforward layers from all pre-trained LLMs into a single MoE module at each attention-FFN block, and merging other layers, such as self-attention layers, by simply averaging their weights (Sukhbaatar et al., 2024). An additional gate network is introduced for each MoE module to perform token-level routing. However, this approach requires extensive fine-tuning of the entire network parameters and the gating network, as there is a significant gap between MoE experts and pre-trained LLMs. Notably, each expert in an MoE tends to become a generalist across all domains due to the load balancing loss, which encourages an even distribution of the workload among experts (Fedus et al., 2022; Jiang et al., 2024). This contrasts with pre-trained expert models, which are typically specialized for specific domains. Compared with MoE, our model-level routing has a significantly lower training cost, as we only need to train a router while keeping the expert models frozen.
Model size reduction analysis. Let $\Psi$ be the model size of an FP16 pre-trained model. When storing $M$ FP16 models, the total model size is $M\Psi$. In contrast, using our method, we store a single base model, a model-level router and $M$ compressed delta models. The total model size is $\Psi+M\tilde{\Psi}+\Phi$, where $\tilde{\Psi}$ represents the size of the compressed delta model and $\Phi$ denotes the storage requirement for the router. Therefore, the compression ratio can be computed by $M\Psi/(\Psi+M\tilde{\Psi}+\Phi)$. For example, when serving 9 LLaMA-13B models, our method achieves a compression ratio of 3.63$\times$. Empirical studies on compression ratios for varying model numbers are shown in Figures 6 and 8.
@@ 5 Experiments
Candidate LLMs. We apply our ME-Switch to two model families, Mistral-7B (Jiang et al., 2023) and LLaMA-2-13B (Touvron et al., 2023). For the Mistral family, we include Dolphin-2.2.1-Mistral-7B  [^1] https://huggingface.co/cognitivecomputations/dolphin-2.2.1-mistral-7b as the instruction expert, Speechless-Code-Mistral-7B  [^2] https://huggingface.co/uukuguy/speechless-code-mistral-7b-v1.0 as the code expert, and MetaMath-Mistral-7B  [^3] https://huggingface.co/meta-math/MetaMath-Mistral-7B as the math expert. For the LLaMA-2-13B family, LLaMA-2-13B-Chat (Touvron et al., 2023) serves as the instruction expert, MetaMath-13B  [^4] https://huggingface.co/meta-math/MetaMath-13B-V1.0 as the math expert, and LLaMA2-Chinese-13B-Chat  [^5] https://huggingface.co/FlagAlpha/Llama2-Chinese-13b-Chat as the Chinese expert. The above models are fine-tuned based on pre-trained backbones. We use Qwen1.5-1.8B-Chat (Bai et al., 2023) as the model-level router.
Training and testing datasets. We collect a diverse set of instruction samples from various open-source datasets, including Alpaca (Taori et al., 2023) for the instruction domain, MetaMathQA (Yu et al., 2023b) for the mathematics domain, Code-74k-ShareGPT  [^6] https://huggingface.co/datasets/ajibawa-2023/Code-74k-ShareGPT for the code domain, and Chinese Alpaca  [^7] https://huggingface.co/datasets/hfl/alpaca_zh_51k for the Chinese domain. To measure the performance of the resulting LLMs, we report accuracy on several benchmarks across different domains: MMLU (Hendrycks et al., 2021a) for the instruction, GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) for the mathematics, HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) for the code, and C-Eval (Huang et al., 2024) and C-MMLU (Li et al., 2023) for the Chinese. We use the WizardCoder toolbox to evaluate on HumanEval and MBPP, and the OpenCompass toolbox (Contributors, 2023) to evaluate on other datasets. For MMLU, we report accuracy based on 5-shot in-context learning. To determine the answer for each question, we assess the perplexity of various response options and select the one with the lowest perplexity. For GSM8K and MATH, we adopt a 4-shot Chain of Thought (CoT) methodology to obtain the final answer following (Wei et al., 2022). For the HumanEval and MBPP datasets, we employ a 0-shot configuration and generate answers using greedy decoding. We assess the functional correctness using the pass@1 metric following (Liu et al., 2024b).
Implementation details. For salient-aware delta compression, we construct a calibration set from each domain-specific dataset and use these sets to compress the delta weights for each respective domain. Each calibration set consists of 1600 randomly sampled sequences, each with a length of 128 tokens. The bitwidth $b$ and the number of FP16 input channels $k$ are set to 2 and 8, respectively. We use the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of $10^{-5}$ and a mini-batch size of 4 for training over 1 epoch. Delta weights compression experiments for the Mistral family are conducted on two NVIDIA A100 80G GPUs, while for the LLaMA-2-13B model, we use four NVIDIA A100 80G GPUs.
@@
ME-Switch w/o Routing 44.57 60.87 64.00 58.04 55.45 70.05 13.20 41.63 40.13 39.91 40.02
ME-Switch 44.51 60.87 64.00 58.04 55.43 69.90 13.14 41.52 40.13 39.84 39.99
5.1 Main Results
To evaluate the efficacy of our proposed model, we apply ME-Switch to the Mistral-7B and LLaMA-13B model families. For results in terms of LLaMA-3-8B model family, please see Section B of the appendix. We include ME-Switch without model-level routing for comparisons. The experimental results, detailed in Table 1, demonstrate that ME-Switch without routing, even with extremely compressed delta weights, achieves performance comparable to that of the respective unquantized expert models across various downstream tasks. For the Mistral-7B family, on MMLU, ME-Switch without routing lags behind the math expert by just 0.22% in mathematical reasoning tasks. Notably, ME-Switch without routing consistently outperforms the code expert in code generation tasks. The performance improvements over uncompressed expert models are primarily attributed to additional training through efficient distillation, which improves the models’ task-specific performance by optimizing the quantization step size. With model-level routing, ME-Switch nearly matches the performance of the ones without it, demonstrating its precise routing for user queries.
Table 2: Performance comparisons between fixed-precision quantization and mixed-precision quantization for MetaMath-Mistral-7B and Speechless-Code-Mistral-7B.
Method Model Size (GB) Mathematical Reasoning (%) $\uparrow$ Code Generation (%) $\uparrow$
GSM8K Math Avg. HumanEval MBPP Avg.
Full-precision 13.48 73.92 20.62 47.27 51.20 60.40 55.80
@@ 5.1 Main Results
Mixed-precision 2.13 73.62 20.48 47.05 51.80 60.70 56.25
Table 3: Effect of different FP16 input channel numbers $k$ for Speechless-Code-Mistral-7B.
Model Model Size (GB) HumanEval MBPP Avg.
FP 13.48 51.20 60.40 55.80
@@
$k=16$ 2.15 49.40 61.90 55.65
$k=32$ 2.19 51.20 61.20 56.20
$k=64$ 2.26 51.20 61.60 56.40
@@ 5.2 Ablation Studies
Fixed-precision quantization vs. Mixed-precision quantization. To validate the effect of mixed-precision quantization, we compress the delta weights of MetaMath-Mistral-7B and Speechless-Code-Mistral-7B using both fixed-precision quantization and our salient-aware mixed-precision quantization. We evaluate their performance on mathematical reasoning and code generation tasks, respectively. The bitwidth $b$ and the number of FP16 input channels $k$ are set to default values as specified in the implementation details. The results in Table 3 indicate that despite retaining a minimal number of FP16 channels, the model size of the mixed-precision model (2.13 GB) is nearly identical compared to fixed-precision model (2.11 GB). However, introducing a small number of FP channels significantly improves performance. For example, our compressed Speechless-Code-Mistral-7B, with a model size reduced by 6.33$\times$, even outperforms the full-precision counterpart by 0.45% in the average accuracy on code generation tasks. This underscores the capability of salient-aware mixed-precision quantization to minimize model size while preserving model performance.
Low-rank adaptation vs. our salient-aware delta compression. In addition to mixed-precision quantization, we can also employ low-rank adaptation (LoRA) to compress delta weights. Specifically, we decompose the delta weights as $\Delta=\mathbf{U}\mathbf{\Sigma}\mathbf{V}$ and approximate delta weights using low-rank approximation $\tilde{\Delta}=\mathbf{A}\mathbf{B}$ where $\mathbf{A}=\tilde{\mathbf{U}}\sqrt{\tilde{\mathbf{\Sigma}}}$ and $\mathbf{B}=\sqrt{\tilde{\mathbf{\Sigma}}}\tilde{\mathbf{V}}$. Subsequently, $\mathbf{A}$ and $\mathbf{B}$ are refined using our efficient distillation mentioned in Section 4.1. To compare the effectiveness of LoRA against mixed-precision quantization, we apply both methods to the delta weights of Dolphin-2.2.1-Mistral-7B and MetaMath-Mistral-7B and evaluate performance on instructional and mathematical reasoning tasks. For LoRA, we set the rank to 512. The results are detailed in Table 5. Our approach with a much smaller model size outperforms LoRA, especially on Math. These results reveal that LoRA cannot accurately approximate delta weights for full fine-tuned models like Dolphin-2.2.1-Mistral-7B and MetaMath-Mistral-7B, given that their delta weights lack low-rank properties, as also shown in Figure A in the appendix.
Effect of different number of FP16 channels. To assess the impact of varying FP16 input channel counts $k$, we compress the delta weights of Speechless-Code-Mistral-7B and evaluate performance on code generation tasks. The bitwidth $b$ is set at 2. From Table 3, our method already achieves lossless performance with $k=8$. Increasing $k$ further yields no significant performance improvements, indicating that performance has plateaued. Therefore, we set $k$ to 8 by default.
Table 4: Performance comparisons between LoRA and our salient-aware delta compression for Dolphin-2.2.1-Mistral-7B and MetaMath-Mistral-7B.
Model Model Size (GB) MMLU (%) $\uparrow$ Mathematical Reasoning (%) $\uparrow$
STEM Hums. Social Other Avg. GSM8K Math Avg.
LoRA 2.99 52.16 68.38 73.15 65.52 63.33 73.62 17.30 45.46
Ours 2.11 53.17 69.09 73.88 65.40 63.95 73.62 20.48 47.05
Table 5: Effect of different bitwidths $b$ for Speechless-Code-Mistral-7B.
Model HumanEval MBPP Avg.
FP 51.20 60.40 55.80
[…]
$b=2$ 51.80 60.70 56.25
$b=4$ 51.80 60.20 56.00
Effect of different quantization bitwidths. To investigate the impact of varying bitwidths $b$, we compress the delta weights of Speechless-Code-Mistral-7B and evaluate the performance on a code generation task. Table 5 shows that increasing $b$ from 1 to 2 significantly improve performance, achieving lossless results. Given that further increasing $b$ lead to negligible performance differences due to saturation, we set $b$ to 2 as the default.
Figure 4: Average accuracy vs. delta weights size across different domains. “Baseline” refers to the fixed-precision quantization baseline. The dashed line indicates the full-precision counterpart.
@@
Performance comparisons with other weight-only quantization methods. To demonstrate the promising performance of our salient-aware delta compression, we include the following weight-only quantization methods: AWQ: we use AWQ (Lin et al., 2024) to rescale input channels of delta weights before quantization to mitigate quantization errors. Random: using our salient-aware delta compression, we randomly select some input channels from the delta weights as important channels. Wanda: leveraging our salient-aware delta compression, we select important input channels of delta weights using the pruning metric from Wanda (Sun et al., 2024). Magnitude: within our salient-aware delta compression framework, we select sensitive input channels of delta weights based on their weight magnitude, following the method proposed by (Dettmers et al., 2022). We also include fixed-precision quantization for comparisons. We applied all methods to compress the delta weights of Dolphin-2.2.1-Mistral-7B, MetaMath-Mistral-7B, and Speechless-Code-Mistral-7B, using bitwidths $b=1$ and $b=2$. The results are shown in Figure 4. The detailed number of different methods can be found at Section C of the appendix. From the results, we observe that AWQ achieves comparable performance to the 1-bit baseline on code domain, highlighting the limitations of rescaling in extremely low-bitwidth quantization. In contrast, keeping the salient delta input channels performs favourably against the rescaling input channel counterpart. Moreover, our salient channel selection demonstrates superior performance than Random, Wanda, and Magnitude metrics across various bitwidths and tasks. For example, for $b=2$, our salient-aware delta compression outperforms Wanda by 1.0% on the average accuracy on code domain, underscoring the effectiveness and superiority of our approach in selecting the salient delta weights.
@@
Model size reduction analysis. To investigate the model size reduction as discussed in Section 4.2, we compare the total storage requirements of full-precision models with those of our compressed models for the Mistral-7B and LLaMA-13B families across varying model counts, as shown in Figures 6 and 8. As the number of models increases, the compression ratios improve substantially. For instance, with nine models, our method achieves a 3.63$\times$ reduction compared to full-precision models for Mistral-7B family. These savings become even more pronounced with larger model sizes, reaching up to a 3.63$\times$ reduction for LLaMA-13B family.
Latency analysis. To assess the latency improvements from delta weights compression, we measured the end-to-end decoding latency of the Mistral-7B model with an input sequence length of 128 on a single NVIDIA A100. Due to page limitations, additional latency decomposition results are provided in Section F of the appendix. Decoding latency is critical, as it typically dominates processing time in LLM operations (Lin et al., 2024; Liu et al., 2023). Our efficient Triton kernel, which enables batched matrix multiplication between multiple compressed weight matrices and high-precision input activations, is compared against the conventional approach of individually processing multiple models. Results depicted in Figure 8 illustrate that while our method may perform slightly slower than the naive approach for a small number of models due to additional dequantization overhead, it provides lower latency as the number of models is greater than 4. More importantly, unlike the naive approach, our method is able to simultaneously serve 16 models on GPUs without running into out-of-memory (OOM) issues, demonstrating better scalability and efficiency in high-load scenarios.
@@ 6 Conclusion and Future Work
Appendix